A particle projected vertically upwards with initial velocity 25m/s. during the 3rd second what is it displacement,distance,speed and velocity. g=10m/s

Grade:12

1 Answers

Apoorva Arora IIT Roorkee
9 years ago
we use the three equations of motion for solving the above question. i.e
v=u+at
s= ut +\frac{1}{2}at^{2}
2as=v^{2}-u^{2}
also, when we put v=0 in the first equation, we get t= 2.5 sec
so 3 sec after means
now u becomes zero and a=+10 also t= 0.5 sec
also in the nth second the displacement is given by
S_{n}=u+\frac{1}{2}a+at
which can be easily calculated.